Graphically derive the saving function if the consumption function is:

C+ a + bY

Note: In your Graph indicate and explain the following positions: Y>C;

Y<C; Y=C


1
Expert's answer
2021-03-08T09:27:55-0500
"solution"

"Saving \ function\ can\ be\ derived from\ the\ consumption\ function\ as\ follows;\\"

"S=f(Y)\\
Y=C+S\\
S=Y-C"


"thus\ savings\ is \ the\ amount \ of income\ not \ spent\ in \ consumption;\\"

"therefore;\\
C=C-cY\\
S=Y-C-cY\\
S=-C+Y(1-c)\\"


"C \ is \ MPC"


"As \ change\ in\ income\ is \ devoted \ either \ to \ a \ change\ in \ consumption\\ or \ a\ change\ in 
 \ saving\ or\ to\ both, \ therefore,\ the \ two \ ratios, \ that\ is;"


"[\frac{∆C}{∆Y }]\ and\ [\frac{ ∆S}{∆Y}]\ should \ add\ up\ to\ 1.\\

MPC + MPS = 1\\


or \ 1 – MPC = MPS \ or\ (1 – c) \ is \ the \ MPS"


"Where\ MPC\ represents\ the \ slope \ of\ consumption \ curve \ and \ MPS\ (1 – c)\\ represents \ the\ slope \ of\ saving \ curve.\\
 

As\ MPS\ is\ positive, \ therefore, \ saving\ is \ an\ increasing \ function\ of\ the\\
 income\ level,\ i.e.,\ saving \ increases\ with\ increase\ in\ income."









"for \ example"

"given;"









"As \ consumption\ is\ directly\ related\ to \ the \ income\ level, \ therefore\\ consumption \ curve\ is\ positively\ sloped. \\C = f(Y)"


"AD \ curve (C +I) \ is\ positively \ sloped.\ The gap \ between\ the\ consumption \\curve\ and\ AD \ curve\ is\ constant, \ that\ is,\ the\ two \ curves\ are\ parallel \\because\ Investment\ is \ constant."


"Saving \ curve\ is \ positively\ sloped \ because\ saving\ is\ directly\ related\\ to\ the\ income\ level. S = f(Y)"







"At\ very \ low\ income\ level\ there\ is\ dissaving. Dissaving\ implies \ that\\ consumption\ is \ greater\ than\ the\ income\ level.\ This\ increased\ consumption\\ is\ met\ from\ past \ savings, \ therefore\ , there\ will \ be\ dissaving \\in\ the\ economy,

e.g.\\ Y = 10\\

C = 15\\

Y = C + S\\"


"S = 10- 15\\

Or, S = -5\ → Dissaving"